# JLQCD collaboration

### v2

@@ Acknowledgements.
We thank T. Cohen, H.-T. Ding, C. Gattringer, L. Glozman, A. Hasenfratz, C.B. Lang, O. Phillipsen, R. Pisarski, S. Prelovsek, for fruitful discussions. Discussions during the long-term workshop, HHIQCD2024, at Yukawa Institute for Theoretical Physics (YITP-T-24-02) were useful to complete this work. We also thank the members of JLQCD collaboration for their encouragement and support. Numerical simulations are performed on IBM System Blue Gene Solution at KEK under a support of its Large Scale Simulation Program (No. 16/17-14), Oakforest-PACS at JCAHPC under a support of the HPCI System Research Projects (Project IDs: hp170061, hp180061, hp190090, hp200086, and hp210104) and Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba (xg17i032 and xg18i023), the supercomputer Fugaku provided by the RIKEN Center for Computational Science (hp200130, hp210165, hp210231, hp220279 and hp230323), Wisteria/BDEC-01 Odyssey at JCAHPC (HPCI: hp220093, hp230070 and hp240050 MCRP: wo22i038), and Polarie and Grand Chariot at Hokkaido University (hp200130). We used Japan Lattice Data Grid (JLDG) for storing a part of the numerical data generated for this work. This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216 and JP18H04484, JP22H01219), and by MEXT as “Priority Issue on Post-K computer” (Elucidation of the Fundamental Laws and Evolution of the Universe) and by Joint Institute for Computational Fundamental Science (JICFuS).
@@ Appendix A Raw correlators
In this appendix, we summarize the results for the raw correlator with the lightest quark mass $m=0.001(2.6\mbox{MeV})$ at the simulated temperatures. Some data having negative values are neglected.

### v4

@@ Acknowledgements.
We thank T. Cohen, H.-T. Ding, C. Gattringer, A. Hasenfratz, C.B. Lang, O. Phillipsen, R. Pisarski, S. Prelovsek, for fruitful discussions. Discussions during the long-term workshop, HHIQCD2024, at Yukawa Institute for Theoretical Physics (YITP-T-24-02) were useful to complete this work. We also thank the members of JLQCD collaboration for their encouragement and support. Numerical simulations are performed on IBM System Blue Gene Solution at KEK under a support of its Large Scale Simulation Program (No. 16/17-14), Oakforest-PACS at JCAHPC under a support of the HPCI System Research Projects (Project IDs: hp170061, hp180061, hp190090, hp200086, and hp210104) and Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba (xg17i032 and xg18i023), the supercomputer Fugaku provided by the RIKEN Center for Computational Science (hp200130, hp210165, hp210231, hp220279 and hp230323), Wisteria/BDEC-01 Odyssey at JCAHPC (HPCI: hp220093, hp230070 and hp240050 MCRP: wo22i038), and Polarie and Grand Chariot at Hokkaido University (hp200130). We used Japan Lattice Data Grid (JLDG) for storing a part of the numerical data generated for this work. This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216 and JP18H04484, JP22H01219), and by MEXT as “Priority Issue on Post-K computer” (Elucidation of the Fundamental Laws and Evolution of the Universe) and by Joint Institute for Computational Fundamental Science (JICFuS).
@@ Appendix A Raw correlators
In this appendix, we summarize the results for the raw correlator with the lightest quark mass $m=0.001(2.6\mbox{MeV})$ at the simulated temperatures. Some data having negative values are neglected. Correlators for two volumes at the lowest temperatures in our study are shown in figure 12, while the largest aspect ratio plots for temperatures above $T_{c}$ are shown in figure 13.
Appendix B Screening Mass Plots for $am>0.0010$
As an additional check of the quark mass dependence on the meson screening mass, we present the same plots as Fig. 4 but with three different quark masses in Fig. 14.
[…]